OpenStudy (anonymous):

convert t = -2(1 + cos theta) to rectangular form

OpenStudy (anonymous):

Help!

OpenStudy (anonymous):

\[t = -2( 1 + \cos \theta) \]

OpenStudy (anonymous):

convert to rectangular form............